OpenStudy (anonymous):

Simplify. 25c +15d/ 50c^2 +30d^2 I am looking for the restrictions for c and d.

OpenStudy (across):

\[\frac{25c+15d}{50c^2+30d^2}=\frac{5(5c+3d)}{10(5c^2+3d^2)}\]

OpenStudy (anonymous):

so, both c and d cannot be zero simultaneously

OpenStudy (anonymous):

that means if c is zero d cannot be zero and if d is zero c cannot be zero

OpenStudy (anonymous):

@SUROJ Would their be a radical restriction. If you set the orignal denominator = to zero, then you would end up taking a square root of a negative fraction and this was an Algebra I (middle school) math problem so they can't do imaginary roots.

OpenStudy (anonymous):

I think you can do that.